OpenStudy (anonymous):

Evaluate 2/5 + (y - 1/5) if y = 1/3.

OpenStudy (anonymous):

your job is to compute \[\frac{2}{5}+(\frac{1}{3}-\frac{1}{5})\]

OpenStudy (anonymous):

I don't know how to solve the 1/3-1/5

OpenStudy (anonymous):

ok to subtract fractions with unlike denominators use \[\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bd}\] either that or a calculator with a fraction key

OpenStudy (anonymous):

so 2/15?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

last job is \(\frac{1}{5}+\frac{2}{15}\) use 15 as your lcd